OpenStudy (anonymous):

10. A cheetah ran 300 feet in 2.92 seconds. What was the cheetah’s average speed in miles per hour?

OpenStudy (anonymous):

@Haseeb96 @texaschic101

OpenStudy (anonymous):

\[\frac{300}{\frac{5280\ 292}{100\ 3600}}=\frac{56250}{803}=70.0498 \]

OpenStudy (anonymous):

70

OpenStudy (anonymous):

\[\frac{300/5280}{\left.\frac{292}{100}\right/3600}\text{=}\frac{56250}{803}\text{=}70.0498 \]Used Mathematica 9 for the initial solution and it simplified when that was not desired. The above is from Mathematica 7.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

You are welcome.